# Function Induction and Task Generalization: An Interpretability Study with Off-by-One Addition

### v1

@@ C.3.1 Logit Lens Analysis
In this section, we apply logit lens [28], a widely-adopted interpretability method, to off-by-one addition. This involves directly computing the logits from intermediate layer representations using the final layer norm and the final unembedding layer.
@@
C.3.2 Activation Patching Analysis
@@
In this section, we apply activation patching [22] to off-by-one addition. We performed this analysis in the early stages of our work to gather initial intuitions and signals for our problem, before transitioning to path patching for a more fine-grained understanding of the model’s internal computation.
@@
Fig. 19(a) visualizes the information flow from in-context examples to the residual stream of the last “ =” token. Additionally, Figure 19(b) highlights several layers, specifically layers 32, 36, and 39 at the last “=” token, and layers 35 and 38 at the answer token $c$ in the in-context examples. This aligns with the FI heads (H36.7, H39.7, H39.12) and PT heads (H35.9, H35.14, H38.6, H38.7, H38.9) identified in §3.2. Figure 19(c) further reveals that MLP layers also play critical roles at certain positions. It is possible that FI heads write the +1 function to the residual stream, with subsequent attention and MLP layers involved in the execution of the +1 function. This hypothesis is inspired by prior observations of how MLP layers in transformer models are involved in arithmetic operations [27, 35]. In this work, we limit our focus to attention heads, deferring further analysis of MLP layers to future work.
@@
C.3.3 Alternative Head Ablation Methods
In the main paper, we “ablate” or “knock out” a head by replacing its output in the $x_{cont}$ forward pass with the corresponding head output in the $x_{base}$ forward pass. This instance-specific ablation approach is adopted to better isolate the +1 function computation in each instance. However, this differs from ablation methods commonly used in interpretability work, such as zero ablation [10] or mean ablation [42].
To demonstrate the consistency of our findings across different ablation settings, we repeated the experiment in Fig. 4(a) using zero ablation and mean ablation. For mean ablation, we averaged head outputs at the final “=” token from 100 standard addition examples. We found that in all ablation settings (zero ablation, mean ablation, and our instance-specific ablation), the contrast accuracy reduced to 0% and the base accuracy increased to 100% after ablation.
Appendix D Function Vector Analysis
Figure 20: Rescaled Effect of H32.6, H25.13 and H32.4. We rescale the effect of these three heads in Fig. 5 to [-0.15, 0.15] to make the patterns more readable. When $x$ is the input, both H25.13 and H32.4 suppresses $x-1$.
Rescaled Effect of Selected Heads.
In Fig. 20, we rescale the effect of H32.6, H25.13, H32.4 in Fig. 5 to visualize their patterns more clearly. H25.13 and H32.4 contribute to the +1 function by suppressing $x-1$. However, the role of H32.6 is unclear from the heatmap.
@@ Appendix D Function Vector Analysis
Our function vector style analysis in §4 primarily focuses on what FI heads write out in off-by-one addition. However, these heads may also assume roles in standard addition. To investigate this, we add the FI head outputs in the $M(.|x_{base})$ to the naive prompt $x_{naive}$, and visualize the effect in Fig. 21. By comparing Fig. 5 and Fig. 21, we observe that most FI heads contribute meaningful but distinct information in standard addition, with H39.12 being an exception given its minimal effect in standard addition. The aggregated effect in the bottom-right panel in Fig. 21 suggests that FI heads collectively suppress $x-1$ and promote $x$ in standard addition.
One possibility is that FI heads reinforce the answer $x$, or double-check it by performing $(x-1)+1$ in standard addition. In contrast, during off-by-one addition, the standard addition answers are first “locked in” after early layers, and the FI heads are repurposed to perform +1. We leave further investigation of this phenomenon to future work.
@@
Previously in Fig. 7, we demonstrated how the effect of H39.7 and H25.13 changes with respect to different offset $k$. In Fig. 22-24 we report the effect of all nine heads when $k=-2,-1,2$. We find that for some heads (e.g., H32.1 and H24.9), their effect of suppressing $x$ remains consistent across different $k$ values. For other heads (e.g., H39.7, H39.12, H25.13), their effect changes accordingly with respect to $k$.
Figure 22: Individual and Overall Effect of FI Heads in Off-by-$k$ Addition, $k=-2$.
Figure 23: Individual and Overall Effect of FI Heads in Off-by-$k$ Addition, $k=-1$.
@@
Appendix E Task Generalization
E.1 Tasks and Data Preparation
@@
[…]
@@
In the early stages of this work, we experimented with multi-character Caesar cipher. To prevent multiple characters from being tokenized as a single unit (e.g., “ew” as one token in Gemma-2’s tokenizer), we used a preceding whitespace (␣) before each character, formatting it as “␣e␣w” so that “␣e” and “␣w” became separate tokens. However, we ultimately focused on one-character Caesar cipher in the experiments because Gemma-2 (9B) has insufficient performance on this task when working with multiple characters. The tokenization-aware formatting was retained. The actual model input will be “␣c -> ␣d” for the example “c -> d”.
@@ E.2 Results
Previously in Fig. 6, we report results on representative cases, e.g., $k=2$ in off-by-$k$ addition, the subject of “high school government and politics” in shifted MMLU. In Fig. 25-28, we report results of the full list of offsets and subjects.
@@ Appendix F Reproducibility
[…]
@@
[…]
Appendix G Limitations
Regarding path patching experiments (§3), the identified circuit has limitations as it does not perfectly satisfy the faithfulness and completeness criteria, even with our best efforts. This challenge arises because achieving simultaneous satisfaction of faithfulness, completeness, and minimality is difficult, as these criteria often regulate each other. Moreover, number tokens are often mapped into a sinusoidal (Fourier) feature space rather than a linear space in language models [27, 54, 55], which further complicates our interpretability analysis. Besides this limitation, our analysis in §4 focused primarily on the identified FI heads and their causal effect on naive prompts. Future research could dive into the details of the previous token heads and the consolidation heads, or further investigate the query-key and output-value circuits of these heads [6].
Regarding task generalization experiments (§5) , our current scope is limited to two-step tasks where the second step involves a shifting-related function. We anticipate that the function induction mechanism could operate on a broader spectrum of functions, which could be investigated in future work. Additionally, the task pairs we investigated are synthetic or algorithmic; further exploration of the behavior of function induction heads on naturally occurring text distributions would be highly valuable.
Appendix H Broader Impact
Our work does not involve model training and thus does not directly introduce new malicious or harmful applications. However, our findings broadly demonstrate that when presented with false information in the prompt, models tend to follow and even generalize such inaccuracies. This observation could offer insights into unintended applications or behaviors of language models, such as the generation of misinformation (e.g., if prompted with the false premise that the Earth is flat, the model might generate a convincing article supporting this claim) or how the model acts overly sycophantically (e.g., if a user believes “1+1=3”, the model might not only endorse this but also extend it to “2+2=5”). It is our hope that the insights from our interpretability analysis can inform efforts to address these critical societal problems.

### v2

@@ D.3.1 Logit Lens Analysis
In this section, we apply logit lens (nostalgebraist, 2020), a widely-adopted interpretability method, to off-by-one addition. This involves directly computing the logits from intermediate layer representations using the final layer norm and the final unembedding layer.
@@
D.3.2 Activation Patching Analysis
@@
In this section, we apply activation patching (Meng et al., 2022) to off-by-one addition. We performed this analysis in the early stages of our work to gather initial intuitions and signals for our problem, before transitioning to path patching for a more fine-grained understanding of the model’s internal computation.
@@
Fig. 20(a) visualizes the information flow from in-context examples to the residual stream of the last “ =” token. Additionally, Figure 20(b) highlights several layers, specifically layers 32, 36, and 39 at the last “=” token, and layers 35 and 38 at the answer tokens $c_{i}$ in the in-context examples. This aligns with the FI heads (H36.7, H39.7, H39.12) and PT heads (H35.9, H35.14, H38.6, H38.7, H38.9) identified in §3.2. Figure 20(c) further reveals that MLP layers also play critical roles at certain positions. It is possible that FI heads write the +1 function to the residual stream, with subsequent attention and MLP layers involved in the execution of the +1 function. This hypothesis is inspired by prior observations of how MLP layers in transformer models are involved in arithmetic operations (Nanda et al., 2023; Stolfo et al., 2023). In this work, we limit our focus to attention heads, deferring further analysis of MLP layers to future work.
@@
D.3.3 Alternative Head Ablation Methods
In the main paper, we “ablate” or “knock out” a head by replacing its output in the $x_{cont}$ forward pass with the corresponding head output in the $x_{base}$ forward pass. This instance-specific ablation approach is adopted to better isolate the +1 function computation in each instance. However, this differs from ablation methods commonly used in interpretability work, such as zero ablation (Halawi et al., 2023) or mean ablation (Wang et al., 2023).
To demonstrate the consistency of our findings across different ablation settings, we repeated the experiment in Fig. 4 using zero ablation and mean ablation. For mean ablation, we averaged head outputs at the final “=” token from 100 standard addition examples. We found that in all ablation settings (zero ablation, mean ablation, and our instance-specific ablation), the contrast accuracy reduced to 0% and the base accuracy increased to 100% after ablation.
Appendix E Circuit Evaluation
Figure 21: Circuit Evaluation.
In §4, we primarily validated the identified circuit using head ablation experiments and causal effect visualizations. Wang et al. (2023) proposed a more rigorous framework for circuit evaluation, based on faithfulness, completeness, and minimality. In the following, we evaluate the identified circuit according to these metrics. Note that we focus on interpreting the “off-by-one” component of the task, rather than the standard addition component. Hence, these circuit evaluation metrics are adapted accordingly to use $F(M,x_{base})$ as a reference point.
The faithfulness metric measures whether a circuit $C$ has a similar performance to the full model $M$, i.e., whether $F(C,x_{cont})$ is close to $F(M,x_{cont})$, with $F(C,x)$ defined earlier in §3.1. We find that $F(M,x_{base})=7.17$, $F(M,x_{cont})=-1.26$, and $F(C,x_{cont})=0.56$, suggesting that $C$ recovers $\frac{7.17-0.56}{7.17-(-1.26)}=78.4\%$ of the performance of $M$.
[…]
Appendix F Circuit Analysis
Due to space limit, we mainly perform circuit analysis on function induction (FI) heads and present the most notable findings in the main paper (§4). In this section, we discuss remaining findings on FI heads in §F.1. We also present additional analysis on consolidation heads in §F.2 and previous token (PT) heads in §F.3.
F.1 Function Induction (FI) Heads
@@
Our function vector style analysis in §4 primarily focuses on what FI heads write out in off-by-one addition. However, these heads may also assume roles in standard addition. To investigate this, we add the FI head outputs in the $M(.|x_{base})$ to the naive prompt $x_{naive}$, and visualize the effect in Fig. 22. By comparing Fig. 5 and Fig. 22, we observe that most FI heads contribute meaningful but distinct information in standard addition, with H39.12 being an exception given its minimal effect in standard addition. The aggregated effect in the bottom-right panel in Fig. 22 suggests that FI heads collectively suppress x1x-1 and promote $x$ in standard addition.
One possibility is that FI heads reinforce the answer $x$, or double-check it by performing (x1)+1(x-1)+1 in standard addition. In contrast, during off-by-one addition, the standard addition answers are first “locked in” after early layers, and the FI heads are repurposed to perform +1. We leave further investigation of this phenomenon to future work.
@@
Previously in Fig. 7, we demonstrated how the effect of H39.7 and H25.13 changes with respect to different offset $k$. In Fig. 23-25 we report the effect of all nine heads when k=2,1,2k=-2,-1,2. We find that for some heads (e.g., H32.1 and H24.9), their effect of suppressing $x$ remains consistent across different $k$ values. For other heads (e.g., H39.7, H39.12, H25.13), their effect changes accordingly with $k$.
Figure 23: Individual and Overall Effect of FI Heads in Off-by-$k$ Addition, k=2k=-2.
Figure 24: Individual and Overall Effect of FI Heads in Off-by-$k$ Addition, k=1k=-1.
@@
F.2 Consolidation Heads
[…]
We observe that three of these heads (H41.4, H40.11, H40.12) are suppressing answers other than $x$, and one head (H41.5) is promoting answers other than $x$. Their aggregated effect leads to promoting $x$ and suppressing $x+1$, which counters the effect brought by FI heads discussed in Fig. 5.
Figure 26: Causal Effect of Consolidation Heads. The aggregated effect of consolidation heads counter the effect of FI heads by promoting $x$ and suppressing $x+1$.
Discussion.
We name these heads as “consolidation heads” based on three observations: (1) they appear in the final two layers; (2) they attend exclusively to the current token and the <bos> token, suggesting that they mainly process information locally at the current token; (3) our causal analysis in Fig. 26 shows that some of them are promoting 3+3=6 and some are promoting 3+3=7 in off-by-one addition, suggesting that they are weighing the two possible outputs collaboratively.
Despite these observations, our understanding on the exact role of these heads, and why they emerge, remain limited. We believe it relates to the broader phenomenon of “negative” behavior in language models, which has been noted as a challenge for current interpretability methods (Sharkey et al., 2025). We hope future work will present a finer-grained interpretation of these heads.
F.3 Previous Token (PT) Heads
Off-by-one Addition Off-by-two Addition
Prompt Setting Base Acc Cont. Acc Base Acc Cont. Acc
Base Prompt 1.00 0.00 1.00 0.00
Contrast Prompt 0.11 0.86 0.04 0.46
Contrast Prompt (Ablate PT Heads) 0.96 0.04 0.95 0.01
Contrast Prompt (Ablate 8 Random Heads) 0.09 0.89 0.05 0.46
Table 5: Head Ablation Experiments with Previous Token Heads.
Head Ablation Experiments.
To validate the role of previous token heads, we first repeat the head ablation experiments in Fig. 4, but we ablated previous token heads instead. We consider both off-by-one addition and off-by-two addition, and use 16-shots in the prompt. We report the results in Table 5. Ablating the previous tokens heads almost completely restores the model’s default behavior. This supports our hypothesis that these heads are critical for inducing the +1 function.
Causal Effect.
[…]
We report the causal effect of these PT head outputs in Fig. 27. Similar to the findings with FI heads, we find that each PT head conveys distinct information. For example, H38.7 promotes $x\pm 1$ and $x\pm 2$, H35.14 promotes $x+1$ and suppresses x1x-1. Collectively, these heads suppress $x$ and promote $x+1$, which aligns with the hypothesized role of PT heads. Unexpectedly, these heads also promotes x1x-1. We attribute this to the task shift from off-by-one addition and the naive prompt, and the token-level averaging operation we employ which may cause loss of information.
Pairs of Heads with Countering Effect.
We notice that two pairs of PT head (H38.6/7 and H31.4/5) demonstrate opposing patterns, and they happen to be in the same group in group query attention. Similarly, two consolidation heads (H41.4/5; Fig. 26) have a similar countering effect. Hence we hypothesize that group query attention may help these heads develop countering or hedging behaviors.
Figure 27: Causal Effect of Previous Token Heads. (*) Effects of H38.9, H35.14 and H29.5 are rescaled to [-0.02, 0.02] to make the patterns more readable.
Appendix G Task Generalization
G.1 Tasks and Data Preparation
@@
We focus on 6 subjects in the MMLU dataset (Hendrycks et al., 2021): high school government and politics, high school US history, US foreign policy, marketing, high school psychology, sociology. We downloaded the MMLU dataset from hendrycks/test. We chose these subjects because Gemma-2 (9B) achieves 90% accuracy with 5 shots on them. For subjects where Gemma-2 (9B) achieves lower accuracies, tracking and analyzing performance on Shift-by-One MMLU becomes challenging, because the model could score points by random guessing. We use 16 shots in the experiments in Fig. 6(b), where the 16 shots combine “validation” and “dev” examples from the MMLU dataset.
@@
In the early stages of this work, we experimented with multi-character Caesar cipher. To prevent multiple characters from being tokenized as a single unit (e.g., “ew” as one token in Gemma-2’s tokenizer), we used a preceding whitespace (␣) before each character, formatting it as “␣e␣w” so that “␣e” and “␣w” became separate tokens. However, we ultimately focused on one-character Caesar cipher in the experiments because Gemma-2 (9B) has insufficient performance on the multi-character version. The tokenization-aware formatting was retained. The actual model input will be “␣c -> ␣d” for the example “c -> d”.
@@ G.2 Results
Previously in Fig. 6, we report results on representative cases, e.g., $k=2$ in off-by-$k$ addition, the subject of “high school government and politics” in shifted MMLU. In Fig. 28-31, we report results of the full list of offsets and subjects.
@@ Appendix H Reproducibility
We primarily use the  transformer-lens (Nanda & Bloom, 2022) library for model inference and interpretability analysis. This library is built on the  transformers (Wolf et al., 2020) library. We have also used the  llm-transparency-tool (Ferrando & Voita, 2024; Tufanov et al., 2024) for early exploration.
@@
Our code can be found at   INK-USC/function-induction.